A compact disc varies it’s angular velocity so that the linear velocity of the location currently scanned is 1.2 m/s.

What is the angular velocity at the inner part of the disc (r = `28mm) at the start?

What is the angular velocity at the outer part of the disc (r = 58mm) at the end?
Answer #1

(a) at r= 28 mm R 1.2 m/s 1 w = r = Zdx10-3 m - 42.86 radis Answer Ibl at r= 58 mm W- 1. v . 1.2 m/s 58x10-3m = 20.69 rad/s A